# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v3

@@ 6 Gradient dynamics
This shows that gradient dynamics with weight decay will pick a lower-order (i.e., simpler) solution, suggests that perfect memorization may not be not favorable in dynamics. The following theorem shows that the dynamics also enjoys asymptotic freedom:
@@
Considering the modified loss of Eqn. 3 with only the first two terms: ~$\tilde{\ell}_{k}:=-2r_{kkk}+\sum_{k_{1}k_{2}}|r_{k_{1}k_{2}k}|^{2}$, if the weights are i.i.d Gaussian and network width $q\rightarrow+\infty$, then $JJ^{*}$ converge to diagonal and the dynamics of MPs is decoupled.
@@ 7 Experiments
Setup. We train the 2-layer MLP on the modular addition task, which is a special case of outcome prediction of Abelian group multiplication. We use Adam optimizer with learning rate $0.01$, MSE loss, and train for $10000$ epochs with weight decays. We tested on $|G|=d\in\{23,71,127\}$. All data are generated synthetically and training/test split is $90\%/10\%$.
Solution Distributions. As shown in Fig. 4, we see order-4 and order-6 solutions in each frequency emerging from well-trained networks on $d=23$. The mixed solution ${\bm{z}}_{F4/6}$ can be clearly observed in a small-scale example (Fig. 7). This is also true for larger $d$ (Fig. 5). Although the model is trained with heavily over-parameterized networks, the final solution order remains constant, which is consistent with Corollary 1. Large weight decay shifts the distribution to the left (i.e., low-order solutions) until model collapses (i.e., all weights become zero), consistent with our Theorem 6 that demonstrates that gradient descent with weight decay favors low-order solutions. Similar conclusions follow for fewer and more overparameterization (Appendix H).
$d$not %non-factorable error (×$\times 10^{-2}$) solution distribution (%) in factorable ones
@@
Table 2: Matches between order-4/6 solutions from gradient descent and those constructed by CoGO. Number of hidden nodes $q=512$ and weight decay is $5\times 10^{-5}$. Around $95\%$ gradient descent solutions are factorable with very small factorization error ($\sim 0.04$ compared to solution norm on the order of $1$). Furthermore, CoGO successfully predicts $\sim 98\%$ of the structure of the empirical solutions, while the remaining $2\%$ are mostly due to insufficient training (i.e., near miss against known theoretical construction). Here ${\bm{z}}_{\xi}$ is defined in Corollary 4, ${\bm{z}}_{\nu}:={\bm{u}}_{\nu}+{\bm{1}}$ is defined in Tbl. 1, and ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is defined in Eqn. 11. The means and their standard deviations are computed over 5 seeds.
@@
Exact match between theoretical construction and empirical solutions. A follow-up question arises: do the empirical solutions match exactly with our constructions? After all, distribution of solution order is a rough metric. For this, we identify all solutions obtained by gradient descent at each frequency, factorize them and compare with theoretical construction up to conjugation/normalization. To find such a factorization, we use exhaustive search (Appendix H).
@@
The answer is yes. Tbl. 2 shows that around $95\%$ of order-4 and order-6 solutions from gradient descent can be factorized into $2\times 2$ and $2\times 3$ and each component matches our theoretical construction in Corollary 2 and 4, with minor variations. Furthermore, when $d$ is large, most of the solutions become order-4, which is consistent with our analysis for mixed solution ${\bm{z}}_{F4/6}$ (Corollary 5) that one order-6 solution in the form of ${\bm{z}}_{\nu=\mathrm{i}}*{\bm{z}}_{\mathrm{syn},\alpha\beta}$ suffices to achieve a global optimizer, with all other frequencies taking order-4s. In fact, for $d=127$, the number of order-6 solution taking the form of ${\bm{z}}_{\nu=\mathrm{i}}*{\bm{z}}_{\mathrm{syn},\alpha\beta}$ is $(d-1)/2\cdot 2.25\%\approx 1.26$, coinciding with the theoretical results.
@@
Implicit Bias of gradient descent. Our construction gives other possible solutions (e.g., ${\bm{z}}_{3\mathrm{c}}*{\bm{z}}_{\mathrm{syn}}$) which are never observed in the gradient solutions. Even for the observed solutions, e.g. ${\bm{z}}_{\nu}*{\bm{z}}_{\mathrm{syn}}$, the distribution of free parameters is highly non-uniform (Fig. 6), showing a strong preference of certain choices. These suggest strong implicit bias in optimization, which we leave for future work.
Figure 6: Distribution of free parameters ($\xi$, $\nu$, $\alpha$ and $\beta$, all with magnitude $1$) in three kinds of gradient descent solutions identified by CoGO. While any value of these parameters makes a global optimizer, gradient descent dynamics has a particular preference in picking them during optimization.
Figure 7: The convergence path of $z_{c\cdot\cdot}$ when training modular addition using Adam optimizer (learning rate $0.05$, weight decay $0.005$). The final solution contains 2 order-6 (${\bm{z}}^{(k)}_{F6}$) and 1 order-4 (${\bm{z}}^{(k)}_{F4}$) solutions. Note that for $z_{c\cdot\cdot}$, unlike Fig. 3, each order-6 solution contains a constant bias term to cancel out the artifacts of order-4 solution (Corollary 5). For each hidden node $j$, once a dominant frequency emerges, others fade away.
@@ 8 Conclusion and future work
In this work, we propose CoGO (Composing Global Optimizers), a theoretical framework that models the algebraic structure of global optimizers when training a 2-layer network on reasoning tasks of Abelian group with $L_{2}$ loss. We find that the global optimizers can be algebraically composed by partial solutions that only fit parts of the loss, using ring operations defined in the weight space of the 2-layer neural networks across different network widths. Under CoGO, we also analyze the training dynamics, show the benefit of over-parameterization, and the inductive bias towards simpler solutions due to topological connectivity between algebraically linked high-order (i.e., involving more hidden nodes) and low-order global optimizers. Finally, we show that the gradient descent solutions exactly match what constructed solutions (e.g. ${\bm{z}}_{F4/6}$ and ${\bm{z}}_{F6}$, see Corollary 5 and Corollary 2).
@@
Develop novel training algorithms. Instead of applying (stochastic) gradient descent to overparameterized networks, CoGO suggests a completely different path: decompose the loss, find the MPs, construct low-order solutions and combine them to achieve the final solutions on the fly using algebraic operations. Such an approach may be more efficient and scalable than gradient descent, due to its factorable nature. Also, our framework works for losses depending on monomial potentials ($L_{2}$ loss is just one example), which opens a new dimension for loss design.
@@
Putting different widths into the same framework. Many existing theoretical works study properties of networks with fixed width. However, CoGO demonstrates that nice mathematical structures emerge when putting networks of different widths together, an interesting direction to consider.
@@
Grokking. When learning modular addition, there exists a phase transition from memorization to generalization during training, known as grokking (Varma et al., 2023; Power et al., 2022), long after the training performance becomes (almost) perfect. Our work may be expanded to a nonuniformly distributed training set to study the dynamics of representation learning on grokking.
@@
Extending to other activations. For other activation than quadratic (e.g., SiLU) with $\sigma(0)=0$, with a Taylor expansion, the same framework may still apply (with higher rank MPs).
Appendix A Decoupling $L_{2}$ Loss (Proof)
@@
Note that the frequency $k$ goes from 00 to $d-1$, where $\phi_{0}\equiv 1$ is the trivial representation (i.e., all $g\in G$ maps to $1$). According to the Fundamental Theorem of Finite Abelian Groups, each finite Abelian group can be decomposed into a direct sum of cyclic groups, and the character function of each cyclic group is exactly (scaled) Fourier bases. Therefore, in Abelian group, $k$ is a multi-dimensional frequency index.  Conrad (2010) shows that $\hat{G}\cong G$ (Theorem 3.13) so each character function $\phi\in\hat{G}$ can also be indexed by $g$ itself. Right now we keep the index $k$.
@@
summing over $\{p_{1},p_{2},p_{3},p_{4},k_{1},k_{2},k_{3},k_{4},k^{\prime}\neq 0,j_{1},j_{2}\}$. Note that since each $p\in\{a,b\}$, there are $162^{4}=$ 16 choices of $(p_{1},p_{2},p_{3},p_{4})$. For notation brevity, we use $(1,3)$ to represent the subset of $p$ that takes the value of $a$ (e.g., $(1,3)$ means that $p_{1}=p_{3}=a$ and $p_{2}=p_{4}=b$). It is clear that for odd assignments such as $(1,2,3)$, since $z_{p0j}=0$, the summation is zero. Then, we only discuss the even cases as follows:
Case 1: $(1,3)$, $(2,4)$, $(1,4)$, $(2,3)$. The 4 cases are identical so we only need to analyze one. We take $(1,3)$ as an example. For $(1,3)$, $p_{1}=p_{3}=a$, $p_{2}=p_{4}=b$ and the only nonzero terms is when $k_{1}+k_{3}=0\mod d$, $k_{2}+k_{4}=0\mod d$, since $\mathbb{E}_{g_{1}}\left[\phi_{k_{1}}(g_{1})\phi_{k_{3}}(g_{1})\right]=\mathbb{I}(k_{1}+k_{3}=0\mod d)$ (and similar in other cases). Then Eqn. 22 becomes:
@@
Case 2: $(1,2)$ and $(3,4)$. The two cases are identical. Take $(1,2)$ as an example. In this case, $p_{1}=p_{2}=a$ and $p_{3}=p_{4}=b$. The only non-zero terms are when $k_{1}+k_{2}=0$, $k_{3}+k_{4}=0$. Then Eqn. 22 becomes:
@@
Case 3: $(1,2,3,4)$ and ()()( ). In this case we have:
@@
Appendix B Semi-ring structure of $\mathcal{Z}$ (Proof)
@@
Appendix C Solution Construction (Proof)
C.1 Construction of Partial Solutions
@@
C.2 Construction of Global Optimizers
@@
Just notice that ${\bm{z}}_{\mathrm{syn}}:=\bm{\rho}({\bm{u}}_{\mathrm{syn}})={\bm{u}}^{2}_{\mathrm{syn}}+{\bm{u}}_{\mathrm{syn}}+{\bm{1}}_{k}$ (superscript $(k)$ are omitted for brevity) makes all MPs in $R_{\mathrm{n}}$, $R_{\mathrm{c}}$ and part of $R_{*}$ (Tbl. 1) equal to 00, except for “aac” and “bbc”, which corresponds to monomial polynomials $r_{akkk}:=\sum_{j}z_{akj}z_{akj}z_{ckj}$ and $r_{bkkk}:=\sum_{j}z_{bkj}z_{bkj}z_{ckj}$. On the other hand, according to Tbl. 1, ${\bm{z}}_{\nu}:={\bm{u}}_{\nu}+{\bm{1}}_{k}$ has $r_{akkk}({\bm{z}}_{\nu})=r_{bkkk}({\bm{z}}_{\nu})=0$. Therefore, using ring homomorphism, we know that for any $r\in R_{\mathrm{n}}\cup R_{\mathrm{c}}\cup R_{*}$, $r({\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu})=0$ and thus $R_{\mathrm{n}}\cup R_{\mathrm{c}}\cup R_{*}$ is the 0-sets.
@@
First, ${\bm{u}}_{\nu=\mathrm{i}}={\bm{u}}_{4\mathrm{c}}$ in Tbl. 1 and thus $\bm{\rho}({\bm{u}}_{\nu=\mathrm{i}})$ has 0-sets $R_{\mathrm{c}}$ and $R_{*}$ except for “$ab\bar{c}$”, which corresponds to MP $r_{k,k,-k}\in R_{\mathrm{c}}$. On the other hand, we have
@@
The second equality is because all $(a,k^{\prime})$ and $(b,k^{\prime})$ entries are 00 except for $k^{\prime}=\pm k_{0}$, and the last equality is because all nonzero entries of ${\bm{z}}^{(k_{0})}_{F6}$ have magnitude $1$.
@@
$[\hat{\bm{z}}^{(k_{0})}_{F6}]_{cmj}=-\frac{4(1+\mathrm{i})}{2\cdot 6}=-\frac{1}{3}(1+\mathrm{i})$ (65)
and $\sum_{k^{\prime}}r_{p,k^{\prime},-k^{\prime},m}({\bm{z}}_{F4/6})=0$ for all $m$. See Fig. 8) for the construction.
@@
C.3 Canonical Forms

### v4

@@ 6 Exploring the solution solution with Gradient dynamics
This shows that gradient dynamics (with weight decay) may pick a lower-order (i.e., simpler) solution. This suggests that gradient dynamics may not favor perfect memorization, which is of high order. We leave it a future work to prove the existence of a path that connects perfect memorization solutions with a lower-order one. The following theorem shows that the dynamics enjoys asymptotic freedom:
@@
Considering the modified loss of Eqn. 3 with only the first two terms: $\tilde{\ell}_{k}:=-2r_{kkk}+\sum_{k_{1}k_{2}}|r_{k_{1}k_{2}k}|^{2}$, if the weights are i.i.d Gaussian and network width $q\rightarrow+\infty$, then $JJ^{*}$ converge to diagonal and the dynamics of SPs is decoupled.
@@
Figure 4: Solution distribution (accumulated over 5 random seeds) over different weight decay regularization for $q=512$, trained with 10k epochs with Adam with learning rate $0.01$ on modular addition (i.e., predicting $a+b\mod d$) with $d\in\{23,71,127\}$. Red dashed lines correspond to order-4/6 solutions.
@@ 7 Experiments
Setup. We train the 2-layer MLP on the modular addition task, which is a special case of outcome prediction of Abelian group multiplication. We use Adam optimizer with learning rate $0.01$, MSE loss, and train for $10000$ epochs with weight decays. We tested on $|G|=d\in\{23,71,127\}$. All data are generated synthetically and training/test split is $90\%/10\%$. Each training with a fixed set of hyperparameter configuration is conducted on NVIDIA V100 for a few minutes.
Solution Distributions. As shown in Fig. 3, we see order-4 and order-6 solutions in each frequency emerging from well-trained networks on $d=23$. The mixed solution ${\bm{z}}_{F4/6}$ can be clearly observed in a small-scale example (Fig. 6). This is also true for larger $d$ (Fig. 4). Although the model is trained with heavily over-parameterized networks, the final solution order remains constant, which is consistent with Corollary 1. Large weight decay shifts the distribution to the left (i.e., low-order solutions) until model collapses (i.e., all weights become zero), consistent with our Theorem 5 that demonstrates that gradient descent with weight decay favors low-order solutions. Similar conclusions follow for fewer and more overparameterization (Appendix I).
$d$ %not %non-factorable error ($\times 10^{-2}$) solution distribution (%) in factorable ones
@@
Table 2: Matches between order-4/6 solutions from gradient descent and those constructed by CoGS. Number of hidden nodes $q=512$ and weight decay is $5\times 10^{-5}$. Around $95\%$ gradient descent solutions are factorable with very small factorization error ($\sim 0.04$ compared to solution norm on the order of $1$). Furthermore, CoGS successfully predicts $\sim 98\%$ of the structure of the empirical solutions, while the remaining $2\%$ are largely due to insufficient training, near miss against known theoretical construction. Here ${\bm{z}}_{\xi}$ is defined in Corollary 4, ${\bm{z}}_{\nu}:={\bm{u}}_{\nu}+{\bm{1}}$ is defined in Tbl. 1, and ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is defined in Eqn. 68. The means/standard deviations are computed over 5 seeds.
@@
Exact match between theoretical construction and empirical solutions. A follow-up question arises: do the empirical solutions match exactly with our constructions? After all, distribution of solution order is a rough metric. For this, we identify all solutions obtained by gradient descent at each frequency, factorize them and compare with theoretical construction up to conjugation/normalization. To find such a factorization, we use exhaustive search (Appendix I).
@@
The answer is yes. Tbl. 2 shows that around $95\%$ of order-4 and order-6 solutions from gradient descent can be factorized into $2\times 2$ and $2\times 3$ and each component matches our theoretical construction in Corollary 2 and 4, with minor variations. Furthermore, when $d$ is large, most of the solutions become order-4, which is consistent with our analysis for mixed solution ${\bm{z}}_{F4/6}$ (Corollary 5) that one order-6 solution in the form of ${\bm{z}}_{\nu=\mathrm{i}}*{\bm{z}}_{\mathrm{syn},\alpha\beta}$ suffices to achieve a global solution, with all other frequencies taking order-4s. In fact, for $d=127$, the number of order-6 solution taking the form of ${\bm{z}}_{\nu=\mathrm{i}}*{\bm{z}}_{\mathrm{syn},\alpha\beta}$ is $(d-1)/2\cdot 2.25\%\approx 1.26$, coinciding with the theoretical results.
@@
Implicit Bias of gradient descent. Our construction gives other possible solutions (e.g., ${\bm{z}}_{3\mathrm{c}}*{\bm{z}}_{\mathrm{syn}}$) which are never observed in the gradient solutions. Even for the observed solutions, e.g. ${\bm{z}}_{\nu}*{\bm{z}}_{\mathrm{syn}}$, the distribution of free parameters is highly non-uniform (see Fig. 12 in Appendix), showing a strong preference of parameters that lead to symmetry. These suggest strong implicit bias in optimization, which we leave for future work.
@@ 8 Conclusion and future work
In this work, we propose CoGS (Composing Global Solutions), a theoretical framework that models the algebraic structure of global solutions when training a 2-layer network on reasoning tasks of Abelian group with $L_{2}$ loss. We find that the global solutions can be algebraically composed by partial solutions that only fit parts of the loss, using ring operations defined in the weight space of the 2-layer neural networks across different network widths. Under CoGS, we also analyze the training dynamics, show the benefit of over-parameterization, and the inductive bias towards simpler solutions due to topological connectivity between algebraically linked high-order (i.e., involving more hidden nodes) and low-order global solutions. Finally, we show that the gradient descent solutions exactly match what constructed solutions (e.g. ${\bm{z}}_{F4/6}$ and ${\bm{z}}_{F6}$, see Corollary 5 and Corollary 2).
@@
Develop novel training algorithms. Instead of applying (stochastic) gradient descent to overparameterized networks, CoGS suggests a completely different path: decompose the loss, find the SPs, construct low-order solutions and combine them to achieve the final solutions on the fly using algebraic operations. Such an approach may be more efficient and scalable than gradient descent, due to its factorable nature. Also, our framework works for losses depending on sum potentials ($L_{2}$ loss is just one example), which opens a new dimension for loss design.
@@
Putting different widths into the same framework. Many existing theoretical works study properties of networks with fixed width. However, CoGS demonstrates that nice mathematical structures emerge when putting networks of different widths together, which is an interesting direction to consider. This is related to dynamically adding/pruning neurons during training Yoon et al. (2017); Yu et al. (2018); Wu et al. (2019).
@@
Grokking. When learning modular addition, there exists a phase transition from memorization to generalization during training, known as grokking (Varma et al., 2023; Power et al., 2022), long after the training performance becomes (almost) perfect. While our work does not directly address grokking, which involves more complicated training dynamics than described in Sec. 6, our framework may be extended to a nonuniformly distributed training set (e.g. some input pairs $(g_{1},g_{2})$ are missing in the training set), in order to study the dynamics of representation learning on grokking.
@@
Extending to other activations and loss functions. For other activations (e.g., SiLU) with $\sigma(0)=0$, with a Taylor expansion, the same framework may still apply, but with higher rank sum potentials (SPs). For other loss functions, we can do a similar Taylor expansion. We leave them for future work.
Appendix A Notation Table
Symbol Description
$\mathbb{C}$ The set of complex numbers. The complex field.
$\mathbb{N}$ The set of natural numbers.
$\mathrm{i}$ The imaginary unit. $\mathrm{i}=\sqrt{-1}$.
$\bar{a}$, $\bar{\bm{a}}$, $\bar{A}$ The complex conjugate of a scalar $a$, a vector $\bm{a}$} or a matrix $A$.
$A^{*}$ The conjugate transpose of matrix $A$. $A^{*}=\bar{A}^{\top}$.
$\mathbb{I}(x)$ The indicator function. $\mathbb{I}(x)=1$ if $x$ is true, otherwise 0.
$G$ The Abelian group to be studied.
$g\in G$ Group element $g$ in $G$.
$g_{1}g_{2}$ Production of group element $g_{1}$ and $g_{2}$ under group multiplication.
$d$ Size of the group $G$. $|G|=d$.
${\bm{e}}_{g}$ One-hot representation of group element $g$. The dimension of ${\bm{e}}_{g}$ is $d$.
$\phi_{k}:G\mapsto\mathbb{C}$ The $k$-th character function of $G$. If $G$ is cyclic and $0\leq g<d$, then $\phi_{k}(g)=e^{\mathrm{i}2\pi kg/d}$.
$\bm{\phi}_{k}\in\mathbb{C}^{d}$ The $k$-th character function in vector form. $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}$.
$P^{\perp}_{1}$ Zero-mean projection matrix $P^{\perp}_{1}\equiv I-\frac{1}{d}{\bm{1}}{\bm{1}}^{\top}$.
$\mathbf{w}_{aj}$, $\mathbf{w}_{bj}$ Fan-in weight vectors for node $j$ in 2-layer networks defined in Eqn. 1.
$\mathbf{w}_{cj}$ The fan-out weight vector for node $j$ in 2-layer networks defined in Eqn. 1.
$\mathcal{Z}_{q}$ Collection of weight (in Fourier space) of all 2-layer networks with $q$ hidden nodes.
$\mathcal{Z}$ Collection of weight (in Fourier space) of all 2-layer networks. $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$.
${\bm{z}}\in\mathcal{Z}$ Weight matrices of one specific instance of 2-layer network.
$\mathrm{ord}({\bm{z}})$ The number of hidden nodes in $\bm{z}$}.
${\bm{z}}_{1}+{\bm{z}}_{2}$, ${\bm{z}}_{1}*{\bm{z}}_{2}$ The ring addition and multiplication (Def. 5).
$r:\mathcal{Z}\mapsto\mathbb{C}$ The sum potential (Def. 6).
$R$ Collection of sum potentials. E.g., $R_{\mathrm{g}}=\{r_{kkk},k\neq 0\}$.
$R_{\mathrm{g}}$, $R_{\mathrm{c}}$, $_{\mathrm{n}}, RR_{*}$ Collections of sum potentials (Lemma 1) that appear in MSE loss function (Eqn. 3).
Table 3: The notation table.
Appendix B Decoupling $L_{2}$ Loss (Proof)
@@
Note that the frequency $k$ goes from 0 to d1d-1, where $\phi_{0}\equiv 1$ is the trivial representation (i.e., all $g\in G$ maps to $1$). According to the Fundamental Theorem of Finite Abelian Groups, each finite Abelian group can be decomposed into a direct sum of cyclic groups, and the character function of each cyclic group is exactly (scaled) Fourier bases. Therefore, in Abelian group, $k$ is a multi-dimensional frequency index.  Conrad (2010) shows that $\hat{G}\cong G$ (Theorem 3.13) so each character function $\phi\in\hat{G}$ can also be indexed by $g$ itself. Right now we keep the index $k$.
@@
summing over $\{p_{1},p_{2},p_{3},p_{4},k_{1},k_{2},k_{3},k_{4},k^{\prime}\neq 0,j_{1},j_{2}\}$. Note that since each $p\in\{a,b\}$, there are $2^{4}=16$ choices of $(p_{1},p_{2},p_{3},p_{4})$. For notation brevity, we use $(1,3)$ to represent the subset of $p$ that takes the value of $a$ (e.g., $(1,3)$ means that $p_{1}=p_{3}=a$ and $p_{2}=p_{4}=b$). It is clear that for odd assignments such as $(1,2,3)$, since $z_{p0j}=0$, the summation is zero. Then, we only discuss the even cases as follows:
Case 1: $(1,3)$, $(2,4)$, $(1,4)$, $(2,3)$. The 4 cases are identical so we only need to analyze one. We take $(1,3)$ as an example. For $(1,3)$, $p_{1}=p_{3}=a$, $p_{2}=p_{4}=b$ and the only nonzero terms is when $k_{1}+k_{3}=0\mod d$, $k_{2}+k_{4}=0\mod d$, since $\mathbb{E}_{g_{1}}\left[\phi_{k_{1}}(g_{1})\phi_{k_{3}}(g_{1})\right]=\mathbb{I}(k_{1}+k_{3}=0\mod d)$ (and similar in other cases). Then Eqn. 21 becomes:
@@
Case 2: $(1,2)$ and $(3,4)$. The two cases are identical. Take $(1,2)$ as an example. In this case, $p_{1}=p_{2}=a$ and $p_{3}=p_{4}=b$. The only non-zero terms are when $k_{1}+k_{2}=0$, $k_{3}+k_{4}=0$. Then Eqn. 21 becomes:
@@
Case 3: $(1,2,3,4)$ and $()$. In this case we have:
@@
Appendix C Semi-ring structure of $\mathcal{Z}$ (Proof)
@@
Appendix D Solution Construction (Proof)
D.1 Construction of Partial Solutions
@@
D.2 Construction of Global Solutions
@@
Just notice that ${\bm{z}}_{\mathrm{syn}}:=\bm{\rho}({\bm{u}}_{\mathrm{syn}})={\bm{u}}^{2}_{\mathrm{syn}}+{\bm{u}}_{\mathrm{syn}}+{\bm{1}}_{k}$ (superscript $(k)$ are omitted for brevity) makes all MPs in $R_{\mathrm{n}}$, $R_{\mathrm{c}}$ and part of $R_{*}$ (Tbl. 1) equal to 0, except for “aac” and “bbc”, which corresponds to monomial polynomials $r_{akkk}:=\sum_{j}z_{akj}z_{akj}z_{ckj}$ and $r_{bkkk}:=\sum_{j}z_{bkj}z_{bkj}z_{ckj}$. On the other hand, according to Tbl. 1, ${\bm{z}}_{\nu}:={\bm{u}}_{\nu}+{\bm{1}}_{k}$ has $r_{akkk}({\bm{z}}_{\nu})=r_{bkkk}({\bm{z}}_{\nu})=0$. Therefore, using ring homomorphism, we know that for any $r\in R_{\mathrm{n}}\cup R_{\mathrm{c}}\cup R_{*}$, $r({\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu})=0$ and thus $R_{\mathrm{n}}\cup R_{\mathrm{c}}\cup R_{*}$ is the 0-sets.
@@
First, ${\bm{u}}_{\nu=\mathrm{i}}={\bm{u}}_{4\mathrm{c}}$ in Tbl. 1 and thus $\bm{\rho}({\bm{u}}_{\nu=\mathrm{i}})$ has 0-sets $R_{\mathrm{c}}$ and $R_{*}$ except for $ab\bar{c}$”, which corresponds to MP $r_{k,k,-k}\in R_{\mathrm{c}}$. On the other hand, we have
@@
The second equality is because all $(a,k^{\prime})$ and $(b,k^{\prime})$ entries are 0 except for $k^{\prime}=\pm k_{0}$, and the last equality is because all nonzero entries of ${\bm{z}}^{(k_{0})}_{F6}$ have magnitude $1$.
@@
$\hat{\bm{z}}^{(k_{0})}_{F6}]_{cmj}=-\frac{4(1+\mathrm{i})}{2\cdot 6}=-\frac{1}{3}(1+\mathrm{i})$ (63)
and $\sum_{k^{\prime}}r_{p,k^{\prime},-k^{\prime},m}({\bm{z}}_{F4/6})=0$ for all $m$. See Fig. 5 for the construction.
@@
Remarks. To construct $\hat{\bm{z}}_{F6}$, in addition to ${\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu=1}$ shown in the main proof, we could use other compositions to achieve the same effects. For example, ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$, where ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is:
$z_{ak\cdot}=[1,\omega_{3}\alpha,\bar{\omega}_{3}\beta],\quad z_{bk\cdot}=[1,\omega_{3}\bar{\alpha},\bar{\omega}_{3}\bar{\beta}],\quad z_{ck\cdot}=[1,\omega_{3},\bar{\omega}_{3}]$ (68)
D.3 Canonical Forms